Grade 12th passInorganic Chemistry

is benzyl carbocation or tertiary carbocation more stable ?

aneesu
9 Years agoGrade 12th pass

2 Answers

Khimraj
9 Years ago
As benzyl carbocation is stable by resonance and tertiary carbocation is stable by Hyperconjugation. And as resonance is more effective than other. So benzyl carbocation is more stable then tertiary carbocation. For isobaric process and n = 1. dQ = 5/2 RdT + RdT = 7/2RdT. And dW = RdT. So ratio of dQ/dT is 7/2.Hope it clears. If still u have doubts then please clarify. If u like answer then please approve the answer.
Khimraj
9 Years ago
Sorry for mixing answer of another question due to some reason. As benzyl carbocation is stable by resonance and tertiary carbocation is stable by Hyperconjugation. And as resonance is more effective than other. So benzyl carbocation is more stable then tertiary carbocation. Hope it clears. If still u have doubts then please clarify. If u like answer then please approve the answer.